The​ risk-free rate is 3​% and you believe that the​ S&P 500's excess return will be 10.6​% over the next year. If you invest in a stock with a beta of 1.1 (and a standard deviation of 30​%),

what is your best guess as to its expected excess return over the next​ year?

The expected excess return over the next year is? (Round to two decimal​ places.)
Answer #1

Market excess return refers to the market return above the risk free rate.
So, here we have (Market return-Risk free rate) over next year=10.6​%
Stock beta=1.1
Excess Return over next year = Beta *(Excess Market Return)
=1.1*10.6​%=0.1166 or 11.66%

So, the expected excess return over the next year is 11.66%